OpenStudy (anonymous):

n-5 over 2=5 help please forever till i'm done or i will get a F

OpenStudy (anonymous):

when u get equation like this, first thing u do is diminish the fraction by multiply both sides with the amount in denominator.

OpenStudy (anonymous):

multiple by sides by 2 to get n-5=10, add 5 on both sides to get n=10+5=15

OpenStudy (anonymous):

let me try it

OpenStudy (anonymous):

SPACENIGHT DO I ADD 5 TO 15 OR 10 I KNOW I ADD 5 TO 5 I'M CONFUSED

OpenStudy (anonymous):

You add 5 to 10 to get 15.

OpenStudy (anonymous):

add 5 on both sides meant: n-5+5 = 10+5 ---> n=15

OpenStudy (anonymous):

OOOH LOL

OpenStudy (anonymous):

HOW ABOUT \[-9+N OVER 4=-7\]

OpenStudy (anonymous):

I MEANT N OVER 4 LOL SORRY

OpenStudy (anonymous):

Add 9 on both sides so you get (n/4)=2. Multiple by sides by 4 to get n=8

OpenStudy (anonymous):

multiple both sides*

OpenStudy (anonymous):

WAT!!!!!!!!!!

OpenStudy (anonymous):

\[-9 + N \over 4\] = 7

OpenStudy (anonymous):

first, multiply EVERYTHING by 4. like so:

OpenStudy (anonymous):

\[4[(-9 + N)/4] = 4[7]\]

OpenStudy (anonymous):

This removes the fraction.

OpenStudy (anonymous):

\[-9+N OVER 4 =-7

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

anyway you still multiply everything by 4.

OpenStudy (anonymous):

this takes away the fraction, leaving you with -9 + N = 28

OpenStudy (anonymous):

I mean -28

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

add 9 to both sides

OpenStudy (anonymous):

N = 19

OpenStudy (anonymous):

(I mean -19)

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

need any other help with this type of problem?